OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (anonymous):

y=ax+b.parallel lines have the same slope. So a=-3x y=-3x+b. Now substitute value of and and y into the new equation to find b. 6=3+b b=3 I hope you cvan do the rest by yourself

OpenStudy (anonymous):

looks like it is y=-3x+3

OpenStudy (anonymous):

ok let me check

OpenStudy (anonymous):

yes you are correct

OpenStudy (anonymous):

can you medal and fan?

OpenStudy (anonymous):

you will have to ask nick. i just came to comment the answer